### v1

@@ Abstract
keywords:
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Development of EndoViT-based federated foundation models for surgical data.
Integrated MAE, adaptive SAM, and SWA for robust federated pretraining.
Evaluated frozen backbone vs. full fine-tuning in downstream tasks.
Full fine-tuning showed superior performance across all surgical tasks.
Validated on scene segmentation, action recognition, and phase recognition with success.
@@ 1 Introduction
The goal of Surgical Data Science (SDS) is to assist surgeons by recording, analyzing, and modeling surgical data (Maier-Hein et al., 2017). AI introduced analysis of laparoscopic video streams has the potential to aid navigation and decision-making, improve surgical outcomes, reduce morbidity, and enable surgeons to learn from recorded procedures (Maier-Hein et al., 2022). To develop such AI-driven assistance systems, robust and powerful models are required to accurately interpret surgical data and provide meaningful insights (Maier-Hein et al., 2022). It has been shown that Foundation Models are demonstrably robust and powerful models (Bommasani et al., 2021). Bommasani et al. (2021) define Foundation Models as follows: “A foundation model is any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks […].” A foundation model for SDS is, for example, the Endoscopic Vision Transformer (EndoViT) (Batić et al., 2024). This model is pretrained on 700,000 surgical images and has demonstrated strong performance on surgical tasks like Surgical Phase Recognition (SPR), Action Triplet Recognition (ATR), and Surgical Scene Segmentation (SSS). Moreover, the pretrained EndoViT model is already providing meaningful representation when the model is fine-tuned on a limited amount of data.
@@
However, a foundation model trained on limited or homogeneous data may fail to generalize across different patient groups, hospitals, or procedures (Jogan et al., 2024). In this context, homogeneous data refers to datasets collected from a single institution, patient population, or surgical procedure, lacking the variability needed for robust generalization. This variability stems from differences in patient anatomy, surgical techniques, equipment, and clinical protocols (Jogan et al., 2024), making the development of foundation models for surgical applications particularly challenging. Moreover, data protection regulations such as the Health Insurance Portability and Accountability Act (HIPAA) and the General Data Protection Regulation (GDPR) impose strict constraints on patient data sharing, making cross-institutional data aggregation legally and organizationally complex. Federated Learning (FL) offers a solution for this predicament by enabling collaborative model training across medical institutions without data sharing, maintaining privacy while leveraging diverse datasets (McMahan et al., 2017; Bommasani et al., 2021). FL faces challenges from heterogeneous data, e.g., label, feature, quantity, and quality skew (Ye et al., 2023), which must be addressed for effective model convergence and generalization. Otherwise federated training fails in comparison to centralized training.
@@
Despite the early success of FL in surgical tasks, as evidenced by Kassem et al. (2023) for SPR, there is a significant gap in the development of foundation models tailored for surgical applications. In contrast, recent work on adaptive federated knowledge injection into medical MRI and CT foundation models (Anonymous, 2024), as well as FL in real-world medical imaging scenarios (Tölle et al., 2025), highlights the potential for scaling such models in other medical domains. However, both works do not yet address surgical data specifically. AI models for SDS trained via FL remain largely task-specific. One reason for this may be the limited availability of large-scale, multicentric surgical datasets. Another important factor is that surgical data tends to be highly heterogeneous and less standardized compared to radiological imaging data. A compelling example of this heterogeneity is the Critical View of Safety (CVS) challenge, which aggregates cholecystectomy videos from 67 surgeons across 57 countries (Eckhoff et al., 2024), illustrating the substantial variability in surgical technique, recording conditions, and clinical environments.
@@
Contributions: By leveraging federated learning, this work introduces a foundation model specifically designed for the surgical domain, enabling collaborative model development without the need for data sharing. To this end, we extend the EndoViT framework for federated training and integrate adaptive Federated Sharpness-Aware Minimization as proposed by Caldarola et al. (2022). Our approach addresses critical challenges in SDS related to data privacy, institutional heterogeneity, and limited data accessibility, achieving performance comparable to centralized training. This demonstrates that federated foundation models can offer a privacy-preserving and robust solution for SDS, with effective collaboration requiring the adaptation of suitable federated learning methods, such as the integration of FedSAM, which accommodates the inherent inter-center data heterogeneity.
@@ 2 Related Work
This work builds on the results of EndoViT by Batić et al. (2024) and Sharpness-Aware Minimization (SAM) of Caldarola et al. (2022). Therefore, we first introduce the key concepts of the EndoViT study and SAM before detailing our adaptation to a FL setting.
@@ 2.1 EndoViT as Centralized Baseline
As first step the EndoViT model, a Vision Transformer (ViT), was pretrained on the Endo700k dataset collection using the Masked Autoencoder (MAE) approach proposed by He et al. (2022). The MAE method is a SSL technique that reconstructs the missing components of the input images. Specifically, the image is initially divided into 256 non-overlapping patches, where 75 % of these patches are masked randomly. Subsequently, the encoder is trained to process the visible patches to generate a visual representation. A lightweight decoder is then employed to reconstruct the masked patches. This MAE approach enables the effective learning of meaningful representations. Mean-squared error (MSE) is used to calculate the reconstruction loss.
@@
For Surgical Scene Segmentation (SSS), the goal is to assign each pixel in an image to a predefined class. A Dense Prediction Transformer (DPT) (Ranftl et al., 2021) is proposed, which uses a Vision Transformer (ViT) backbone instead of traditional convolutional layers, paired with a convolutional decoder. This architecture, which avoids explicit down-sampling, preserves higher resolution and enables accurate, globally coherent predictions. CholecSeg8k (Hong et al., 2020) is used as dataset for this downstream task. As metric, the Intersection over Union (IoU) is reported for all SSS experiments.
Action Triplet Recognition (ATR) involves identifying the relevant tool, action, and target organ from an image presented as a triplet <instrument, verb, target >. For this task, the authors recommended a model in EndoViT with a feature extraction backbone and a linear head. The evaluation is conducted exclusively on the fifth fold of the ATR split of the CholecT45 dataset (Nwoye and Padoy, 2022) to ensure a fair and representative performance assessment. This particular fold is selected due to its optimal balance in class distribution (Batić et al., 2024; Nwoye and Padoy, 2022). For ATR the mean Average Precision (mAP) is reported.
In Surgical Phase Recognition (SPR), the goal is to classify each video frame into predefined surgical phases. In the EndoViT study, the TeCNO model (Czempiel et al., 2020) is used, which combines Multi-Stage Temporal Convolutional Networks (MS-TCNs) with the EndoViT encoder as a backbone. The large temporal receptive field of TeCNO is particularly advantageous for efficient training and inference on long, unedited surgical videos. Cholec80 (Twinanda et al., 2016) is the dataset used for this downstream task. The metric used for SPR is accuracy (acc).
@@ 2.2 Adaptive Federated Sharpness-Aware Minimization
Caldarola et al. (2022) address the challenges posed by degraded performance and limited generalization in FL settings where data is heterogeneous. The authors argue that this problem is linked to the sharpness of the loss landscape, where sharper minima correlate with poorer generalization. They propose two strategies to mitigate this.
@@
Secondly, they employ Stochastic Weight Averaging (SWA) on the server to aggregate client models. SWA involves averaging weights over multiple training epochs, promoting convergence to flatter minima. As demonstrated by Izmailov et al. (2018), this process tends to identify broader minima in the loss landscape, which are associated with superior generalization. This is due to the averaging of weights over time, which helps to mitigate sharp minima and consequently leads to a global model that exhibits superior generalization to unseen data.
@@ 3.1 Federated EndoViT
Figure 1: Federated EndoViT Framework: A two-stage approach for training a surgical Foundation Model using Federated Learning. (I) The model undergoes federated self-supervised pretraining with a Masked Autoencoder strategy, where 75% of the 256 image patches are masked and reconstructed. Adaptive Federated Sharpness-Aware Minimization is applied on the client side, while Stochastic Weight Averaging is used on the server side to enhance generalization across centers. (II) The pretrained EndoViT encoder serves as a Foundation Model, enabling fine-tuning and feature extraction for diverse surgical downstream tasks. This approach builds upon the work of Batić et al. (2024).
@@
A prevailing trend in Foundation Models is to maintain the frozen state of the model (Rabbani and Bartoli, 2024; Fu et al., 2024; Chen et al., 2024). This practice is driven by numerous factors. One such reason is the potential to reduce training costs and accelerate convergence. Additionally, a pretrained backbone captures general surgical features. Freezing the backbone ensures that task-specific training does not overwrite essential features and avoids overfitting, thereby maintaining the backbone’s robustness and generalizability. To this end, we have conducted a series of ablation experiments in which the backbone is maintained in a frozen state. These experiments are designed to ascertain the potential for enhanced efficiency in the utilization of both FL-EndoViT and CEN-EndoViT.
In addition to reporting acc for SPR, we also provide the F1-score. As discussed by Funke et al. (2023), relying solely on accuracy can be misleading when evaluating classification performance on imbalanced datasets like Cholec80, as it doesn’t account for unequal class distributions and may lead to overly optimistic assessments.
@@ 3.2 Implementation Details
The authors of EndoViT have combined nine publicly available surgical datasets into the Endo700k dataset collection (Batić et al., 2024) with approximately 700,000 images. Each dataset consists of raw minimally invasive surgery (MIS) images from a range of different procedures such as prostatectomy, cholecystectomy, gastrectomy, proctocolectomy, rectal resection and sigmoid resection performed both laparoscopically and robotically.
@@
We implemented our training procedures using the Flower FL Framework (Beutel et al., 2020) in simulation mode, with PyTorch 1.13.0. In our federated setup, each of the nine Endo700k datasets serves as a distinct center. Each center stores its best local model, and the server maintains the best averaged global model. Pretraining in simulation mode was carried out in parallel on four NVIDIA V100 GPUs that do not share memory, while the centralized baseline was retrained on a single NVIDIA V100 GPU. Centralized fine-tuning for SSS, ATR, and SPR was performed on a NVIDIA V100 GPU. Hyperparameters were adopted from the EndoViT Git (Batić et al., 2024) repository to ensure comparability. Our code is available under following link.
@@ 4.1 Federated Pretraining Ablation Study
Our results, shown in Table 1, indicate that the federated MAE approach alone underperforms across all three EndoViT models, each pretrained for a different downstream task in the federated setting. A graphical representation of these results is available in the appendix Section A. Integrating adaptive FedSAM reduces reconstruction error and increases the number of accurately reconstructed patches across different thresholds. However, while it improves upon standard federated training, it does not fully match the performance of the centralized model, though it remains competitive.
While the SWA-enhanced model does not surpass the standard model in validation metrics, it remains valuable due to its potential for improved generalization. Caldarola et al. (2022) suggests that SWA smoothens the optimization landscape by averaging model weights over multiple epochs, allowing convergence toward flatter minima and improving robustness to data variations. This effect is particularly relevant in FL, where heterogeneous data poses a challenge. Additionally, SWA improves stability and consistency across datasets. Previous studies indicate that combining adaptive FedSAM with SWA enhances data augmentation, leading to flatter minima and potentially better generalization than adaptive FedSAM alone. However, our results do not provide conclusive evidence of superior generalization in this specific setup.
@@ 4.2 Downstream Tasks
All experiments are conducted in a centralized environment. To ensure robustness and reliability, we perform three independent technical replications. For few-shot experiments, we use different sets of videos across runs to further assess model consistency. The results reported in this paper represent the average over all technical replications, while the detailed results for each individual replication are provided in Appendix B.1, along with a breakdown of the video assignments across settings.
This experimental setup allows for a direct comparison between the performance of the FL-EndoViT and its centralized counterpart CEN-EndoViT. A graphical representation of the results is provided in Appendix B.
@@ 4.2.2 Fine-tuning Experiment
Fine-Tuned on Resolution CEN-EndoViT FL-EndoViT
@@ B.3 Frozen Backbone Experiments
[…]
Acknowledgements
Funded or Co-funded by the European Union through NEARDATA under grant agreement ID 101092644. Funded or Co-funded by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden. The authors acknowledge the financial support by the Federal Ministry of Education and Research of Germany in the programme of “Souverän. Digital. Vernetzt.”. Joint project 6G-life, project identification number: 16KISK001K.
Disclosure of Interest
Declaration of generative AI and AI-assisted technologies in the writing process.
During the preparation of this work, the author(s) used ChatGPT and DeepL Write in order to improve the readability and language of the manuscript. After using these tools/services, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.

### v2

@@ Abstract
1 Graphical Abstract
@@ 2 Introduction
The goal of Surgical Data Science (SDS) is to assist surgeons by recording, analyzing, and modeling surgical data [1]. AI introduced analysis of laparoscopic video streams has the potential to aid navigation and decision-making, improve surgical outcomes, reduce morbidity, and enable surgeons to learn from recorded procedures [2]. To develop such AI-driven assistance systems, robust and powerful models are required to accurately interpret surgical data and provide meaningful insights [2]. It has been shown that Foundation Models are demonstrably robust and powerful models [3]. Bommasani et al. [3] define Foundation Models as follows: “A foundation model is any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks […].” A foundation model for SDS is, for example, the Endoscopic Vision Transformer (EndoViT) [4]. This model is pretrained on 700,000 surgical images and has demonstrated strong performance on surgical tasks like Surgical Phase Recognition (SPR), Action Triplet Recognition (ATR), and Surgical Scene Segmentation (SSS). Moreover, the pretrained EndoViT model is already providing meaningful representation when the model is fine-tuned on a limited amount of data.
@@
However, a foundation model trained on limited or homogeneous data may fail to generalize across different patient groups, hospitals, or procedures [5]. In this context, homogeneous data refers to datasets collected from a single institution, patient population, or surgical procedure, lacking the variability needed for robust generalization. This variability stems from differences in patient anatomy, surgical techniques, equipment, and clinical protocols [5], making the development of foundation models for surgical applications particularly challenging. Moreover, data protection regulations such as the Health Insurance Portability and Accountability Act (HIPAA) and the General Data Protection Regulation (GDPR) impose strict constraints on patient data sharing, making cross-institutional data aggregation legally and organizationally complex. Federated Learning (FL) offers a solution for this predicament by enabling collaborative model training across medical institutions without data sharing, maintaining privacy while leveraging diverse datasets [6, 3]. FL faces challenges from heterogeneous data, e.g., label, feature, quantity, and quality skew [7], which must be addressed for effective model convergence and generalization. Otherwise federated training fails in comparison to centralized training.
@@
Despite the early success of FL in surgical tasks, as evidenced by Kassem et al. [8] for SPR, there is a significant gap in the development of foundation models tailored for surgical applications. In contrast, recent work on adaptive federated knowledge injection into medical MRI and CT foundation models [9], as well as FL in real-world medical imaging scenarios [10], highlights the potential for scaling such models in other medical domains. However, both works do not yet address surgical data specifically. AI models for SDS trained via FL remain largely task-specific. One reason for this may be the limited availability of large-scale, multicentric surgical datasets. Another important factor is that surgical data tends to be highly heterogeneous and less standardized compared to radiological imaging data. A compelling example of this heterogeneity is the Critical View of Safety (CVS) challenge, which aggregates cholecystectomy videos from 67 surgeons across 57 countries [11], illustrating the substantial variability in surgical technique, recording conditions, and clinical environments.
@@
Contributions: By leveraging federated learning, this work introduces a foundation model specifically designed for the surgical domain, enabling collaborative model development without the need for data sharing. To this end, we extend the EndoViT framework for federated training and integrate adaptive Federated Sharpness-Aware Minimization as proposed by Caldarola et al. [12]. Our approach addresses critical challenges in SDS related to data privacy, institutional heterogeneity, and limited data accessibility, achieving performance comparable to centralized training. This demonstrates that federated foundation models can offer a privacy-preserving and robust solution for SDS, with effective collaboration requiring the adaptation of suitable federated learning methods, such as the integration of FedSAM, which accommodates the inherent inter-center data heterogeneity.
@@ 3 Related Work
This work builds on the results of EndoViT by Batić et al. [4] and Sharpness-Aware Minimization (SAM) of Caldarola et al. [12]. Therefore, we first introduce the key concepts of the EndoViT study and SAM before detailing our adaptation to a FL setting.
@@ 3.1 EndoViT as Centralized Baseline
As first step the EndoViT model, a Vision Transformer (ViT), was pretrained on the Endo700k dataset collection using the Masked Autoencoder (MAE) approach proposed by He et al. [13]. The MAE method is a SSL technique that reconstructs the missing components of the input images. Specifically, the image is initially divided into 256 non-overlapping patches, where 75 % of these patches are masked randomly. Subsequently, the encoder is trained to process the visible patches to generate a visual representation. A lightweight decoder is then employed to reconstruct the masked patches. This MAE approach enables the effective learning of meaningful representations. Mean-squared error (MSE) is used to calculate the reconstruction loss.
@@
For Surgical Scene Segmentation (SSS), the goal is to assign each pixel in an image to a predefined class. A Dense Prediction Transformer (DPT) [14] is proposed, which uses a Vision Transformer (ViT) backbone instead of traditional convolutional layers, paired with a convolutional decoder. This architecture, which avoids explicit down-sampling, preserves higher resolution and enables accurate, globally coherent predictions. CholecSeg8k [15] is used as dataset for this downstream task. As metric, the Intersection over Union (IoU) is reported for all SSS experiments.
Action Triplet Recognition (ATR) involves identifying the relevant tool, action, and target organ from an image presented as a triplet <instrument, verb, target >. For this task, the authors recommended a model in EndoViT with a feature extraction backbone and a linear head. The evaluation is conducted exclusively on the fifth fold of the ATR split of the CholecT45 dataset [16] to ensure a fair and representative performance assessment. This particular fold is selected due to its optimal balance in class distribution [4, 16]. For ATR the mean Average Precision (mAP) is reported.
In Surgical Phase Recognition (SPR), the goal is to classify each video frame into predefined surgical phases. In the EndoViT study, the TeCNO model [17] is used, which combines Multi-Stage Temporal Convolutional Networks (MS-TCNs) with the EndoViT encoder as a backbone. The large temporal receptive field of TeCNO is particularly advantageous for efficient training and inference on long, unedited surgical videos. Cholec80 [18] is the dataset used for this downstream task. The metric used for SPR is accuracy (acc).
@@ 3.2 Adaptive Federated Sharpness-Aware Minimization
Caldarola et al. [12] address the challenges posed by degraded performance and limited generalization in FL settings where data is heterogeneous. The authors argue that this problem is linked to the sharpness of the loss landscape, where sharper minima correlate with poorer generalization. They propose two strategies to mitigate this.
@@
Secondly, they employ Stochastic Weight Averaging (SWA) on the server to aggregate client models. SWA involves averaging weights over multiple training epochs, promoting convergence to flatter minima. As demonstrated by Izmailov et al. [19], this process tends to identify broader minima in the loss landscape, which are associated with superior generalization. This is due to the averaging of weights over time, which helps to mitigate sharp minima and consequently leads to a global model that exhibits superior generalization to unseen data.
@@ 4.1 Federated EndoViT
Figure 1: Federated EndoViT Framework: A two-stage approach for training a surgical Foundation Model using Federated Learning. (I) The model undergoes federated self-supervised pretraining with a Masked Autoencoder strategy, where 75% of the 256 image patches are masked and reconstructed. Adaptive Federated Sharpness-Aware Minimization is applied on the client side, while Stochastic Weight Averaging is used on the server side to enhance generalization across centers. (II) The pretrained EndoViT encoder serves as a Foundation Model, enabling fine-tuning and feature extraction for diverse surgical downstream tasks. This approach builds upon the work of Batić et al. [4].
@@
A prevailing trend in Foundation Models is to maintain the frozen state of the model [20, 21, 22]. This practice is driven by numerous factors. One such reason is the potential to reduce training costs and accelerate convergence. Additionally, a pretrained backbone captures general surgical features. Freezing the backbone ensures that task-specific training does not overwrite essential features and avoids overfitting, thereby maintaining the backbone’s robustness and generalizability. To this end, we have conducted a series of ablation experiments in which the backbone is maintained in a frozen state. These experiments are designed to ascertain the potential for enhanced efficiency in the utilization of both FL-EndoViT and CEN-EndoViT.
In addition to reporting acc for SPR, we also provide the F1-score. As discussed by Funke et al. [23], relying solely on accuracy can be misleading when evaluating classification performance on imbalanced datasets like Cholec80, as it doesn’t account for unequal class distributions and may lead to overly optimistic assessments.
@@ 4.2 Implementation Details
The authors of EndoViT have combined nine publicly available surgical datasets into the Endo700k dataset collection [4] with approximately 700,000 images. Each dataset consists of raw minimally invasive surgery (MIS) images from a range of different procedures such as prostatectomy, cholecystectomy, gastrectomy, proctocolectomy, rectal resection and sigmoid resection performed both laparoscopically and robotically.
@@
We implemented our training procedures using the Flower FL Framework [24] in simulation mode, with PyTorch 1.13.0. In our federated setup, each of the nine Endo700k datasets serves as a distinct center. Each center stores its best local model, and the server maintains the best averaged global model. Pretraining in simulation mode was carried out in parallel on four NVIDIA V100 GPUs that do not share memory, while the centralized baseline was retrained on a single NVIDIA V100 GPU. Centralized fine-tuning for SSS, ATR, and SPR was performed on a NVIDIA V100 GPU. Hyperparameters were adopted from the EndoViT Git [4] repository to ensure comparability. Our code is available under following link.
@@ 5.1 Federated Pretraining Ablation Study
Our results, shown in Table 1, indicate that the federated MAE approach alone underperforms across all three EndoViT models, each pretrained for a different downstream task in the federated setting. Integrating adaptive FedSAM reduces reconstruction error and increases the number of accurately reconstructed patches across different thresholds. However, while it improves upon standard federated training, it does not fully match the performance of the centralized model, though it remains competitive.
While the SWA-enhanced model does not surpass the standard model in validation metrics, it remains valuable due to its potential for improved generalization. Caldarola et al. [12] suggests that SWA smoothens the optimization landscape by averaging model weights over multiple epochs, allowing convergence toward flatter minima and improving robustness to data variations. This effect is particularly relevant in FL, where heterogeneous data poses a challenge. Additionally, SWA improves stability and consistency across datasets. Previous studies indicate that combining adaptive FedSAM with SWA enhances data augmentation, leading to flatter minima and potentially better generalization than adaptive FedSAM alone. However, our results do not provide conclusive evidence of superior generalization in this specific setup.
@@ 5.2 Downstream Tasks
All experiments are conducted in a centralized environment. To ensure robustness and reliability, we perform three independent technical replications. For few-shot experiments, we use different sets of videos across runs to further assess model consistency. The results reported in this paper represent the average over all technical replications.
This experimental setup allows for a direct comparison between the performance of the FL-EndoViT and its centralized counterpart CEN-EndoViT.
@@ 5.2.2 Fine-tuning Experiment
Fine-Tuned on Res. CEN-EndoViT FL-EndoViT
@@ 7 Conclusion
Disclosure of Interest
Acknowledgements
This work was co-funded by the European Union through NEARDATA under grant agreement ID 101092644, the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden, and the Federal Ministry of Education and Research of Germany in the programme of “Souverän. Digital. Vernetzt.”, a joint project 6G-life with the project identification number 16KISK001K.